OpenStudy (anonymous):

Write the equation of the line that passes through (-1, 5) with a slope of 2 in standard form

OpenStudy (anonymous):

A. -2x - y = -7 B. 2x - y = -7 C. 2x - y = 11 D. -2x + y = -11

OpenStudy (anonymous):

Can someone actually work through the problem with me and help me learn it?

OpenStudy (ragingsquirrel):

Give me a minute to type up a reply.

OpenStudy (anonymous):

okay, I also one thats about the same but in point slope form if you could help me on that also!

OpenStudy (anonymous):

thank you so much by the way!

OpenStudy (ragingsquirrel):

Let's do point slope first. I don't know if you'll understand the first question without it.

OpenStudy (jennithemeani):

First thing you have to do is find for B to find for B you need to use point slop form y-y1=m(x-x1) Fill it in and you get y-5= 2(x+1)

OpenStudy (danjs):

Are you familiar with the three basic form for a line?

OpenStudy (jennithemeani):

oops no wait sorry about that doing too many things at once to find b you have to put in y=mx+b which would be 5=2(-1)+b

OpenStudy (jennithemeani):

Which you would make 5=-2+B b=7

OpenStudy (danjs):

k, if you have questions later ill proly see it

OpenStudy (danjs):

What are you doing needing to find the y-intercept. They want the standard form of a line. You have the point-slope form copied down right, all you have to do is rearrange that to ax + by = c standard form

OpenStudy (danjs):

point-slope equation from givens y-y1 = m(x-x1) y - 5 = 2(x - (-1)) move things around to standard form... y - 5 = 2x + 2 -2x + y = 2 + 5 -2x + y = 7 which isn't a choice, multiply everything by -1 2x - y = -7 that's it